The area of a circle that measures 1000m in diameter,is
Here are the four answers given to choose from:
A. 1,000,000m^2
B. 53,930m^2
C. 3,141,592m^2
D. 785,000m^2

Found 2 solutions by jim_thompson5910, jerryguo41:
Answer by jim_thompson5910(35256)

A = pi*r^2

A = pi*(500)^2 .... Note: the radius is half the diameter

A = pi*250000

A = 250000pi

A = 250000(3.14)

A = 785000

So the answer is choice D. 785,000m^2
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------

Answer by jerryguo41(197)

Area of Circle is πr^2
If diameter of circle is 1000m then radius is 500m or one half. Then plug into equation.
A=πr^2
A=π500^2
A=250000π or approx. 785,000m^2